OpenStudy (anonymous):

How many arrangements are possible using the letters in the word FUZZY if each letter “Z” is distinctly different than the other? How many arrangements are possible if the letter “Z” is interchangeable with the other? help i dont know how to do this .can someone please explain step by step how to do this problem to me

jimthompson5910 (jim_thompson5910):

You have 6 letters, so there are 6! = 720 different ways to arrange them if the Z's are distinguishable

jimthompson5910 (jim_thompson5910):

If they are not distinguishable, then 720 is too high of a count and you have to divide by 2! = 2 to correct for this over-counting

OpenStudy (anonymous):

i just have a ? how did you get that there are 6 letters?

jimthompson5910 (jim_thompson5910):

FUZZY has 6 letters

jimthompson5910 (jim_thompson5910):

oh wait...lol miscounted...it should be 5

jimthompson5910 (jim_thompson5910):

looked at it too quickly

jimthompson5910 (jim_thompson5910):

so it would be 5! = 120 and 120/2 = 60

OpenStudy (anonymous):

thank you thats what i thought lol but i didnt know if there was more to it ......

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

could you help me with another ?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

the police department is keeping track of distracted drivers and accidents. they have found that if a driver is distracted, the driver has a 30% chance of being in an accident. if the driver is not distracted, the driver has a 2% chance of being in an accident. the probability of a driver being distracted is 10%. if needed, create a tree diagram. what is the probability a driver will be in an accident? explain. 2. what is the probability that a driver who was in an accident was distracted? explain.

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

ok so we'll have this tree diagram 

jimthompson5910 (jim_thompson5910):

if you multiply out each branch, you'll get this 

jimthompson5910 (jim_thompson5910):

so the probability of getting into an accident is 0.03 + 0.018 = 0.048 this is if you're distracted or not distracted

jimthompson5910 (jim_thompson5910):

So we want to find P(distracted|accident) ------------------------------------------------------- P(distracted|accident) = P(distracted and accident)/P(accident) P(distracted|accident) = 0.03/0.048 P(distracted|accident) = 0.625 which is a 62.5% chance

OpenStudy (anonymous):

thank you very much now i get it.

jimthompson5910 (jim_thompson5910):

yw